An engine pumps a liquid of density 'd' continuously through a pipe of area of cross-section A.If the speed with which the liquid passes through  a pipe is v,then the rate of liquid flow is


a.(Adv³)/2


b.(1/2)Adv


c.(Adv²)/2


d.Adv²

Grade:12th Pass

1 Answers

Suresh
6 years ago
Rate of liquid flow =(1/2*m*v^2)/tRate of flow =m/tRate of flow=dV/tRate of flow=dAL/t1/2* d*A*v *v^21/2 *dAv^3